fe302+4h2=3fe+4h2o when 16.0g of fe3o2 react with excess hydrogen 5.9g fe prduced what is the yeild
1
Expert's answer
2020-02-26T05:01:15-0500

Fe3O2 + 4H2 = 3Fe + 4H2O

16 g x g

(56*3+16*2)g 3*56 g

x = 13.44 g

yeild = 5.9 / 13.44 * 100% = 43.9%